OpenStudy (anonymous):

2. Alice Hall, who loves to cook, makes an apple cake (serves 6) for her family. The recipe calls for 2 1/2 pounds of apples, 2 1/4 cups of flour, 1/5 cup of margarine, 1 1/4 cups of sugar and 4 eggs. Since guests are coming, she would like to make this cake so it will serve 24. How many pounds of apples should she use?

OpenStudy (angela210793):

6----->2 1/2 24--->x x=24*2 1/2/6=4*2 1/2=4*5/2=10 pounds of apples

OpenStudy (anonymous):

The number of servings is 4 times greater. Imagine that she is going to make the cake 4 times. Each time she needs 2 1/2 lbs of apples. so 4 * (2 1/2) = 10 lbs.